# arXiv:2603.01932v1[cs.CV] 02 Mar 2026

### v1

@@ 6 Conclusions
In this paper, we presented BAWSeg, a four-year UAV multispectral benchmark for barley weed segmentation collected over two commercial paddocks in Western Australia. BAWSeg provides radiometrically calibrated five-band reflectance orthomosaics (blue, green, red, red edge, and near-infrared), derived vegetation-index maps, and dense pixel annotations for crop, weed, and other classes, together with leakage-free spatial block splits and deployment-oriented evaluation protocols. Building on this benchmark, we proposed VISA, a two-stream segmentation architecture that decouples radiance-based learning from vegetation-index reasoning and fuses both cues at native resolution to improve robustness under mixed canopy conditions. Across the BAWSeg evaluation suite, VISA achieves a within-plot mIoU of 0.756 with weed IoU of 0.635 and overall accuracy of 0.946 using 22.8M parameters, and it retains mIoU of 0.712 under cross-plot transfer and 0.692 under cross-year testing.
Beyond accuracy gains, our results show that explicitly modeling vegetation-index structure complements calibrated radiance features and improves weed sensitivity under challenging field variability. The ablation study supports this conclusion: removing the index branch yields a 0.025 mIoU decrease, indicating that index-driven context contributes materially to final predictions. BAWSeg and VISA together offer a reproducible foundation for benchmarking multispectral crop–weed segmentation under realistic spatial and temporal shifts, and for developing models that better generalize across seasons and paddocks. To facilitate follow-up research and practical adoption, we will release the BAWSeg dataset, the VISA implementation, and trained model weights upon publication.
@@
Not applicable.
@@
Not applicable.
[…]
Not applicable.
[…]
The BAWSeg dataset, the VISA implementation, and trained model weights will be made publicly available upon publication of this article. The public repository links and persistent identifiers (DOI), as well as detailed usage documentation, will be provided in the final published version.

### v2

@@ 6 Conclusions
In this paper, we presented BAWSeg, a four-year UAV multispectral benchmark for barley weed segmentation collected over two commercial paddocks in Western Australia. BAWSeg provides radiometrically calibrated five-band reflectance orthomosaics (blue, green, red, red edge, and near-infrared), derived vegetation-index maps, and dense pixel annotations for crop, weed, and other classes, together with leakage-free spatial block splits and deployment-oriented evaluation protocols. Building on this benchmark, we proposed VISA, a two-stream segmentation architecture that decouples radiance-based learning from vegetation-index reasoning and fuses both cues at native resolution to improve robustness under mixed-canopy conditions. Across the BAWSeg evaluation suite, VISA achieves a within-plot mIoU of 0.756 with a weed IoU of 0.635 and overall accuracy of 0.946 using 22.8 M parameters, and it retains an mIoU of 0.712 under cross-plot transfer and 0.692 under cross-year testing.
Beyond accuracy gains, our results show that explicitly modelling vegetation-index structure complements calibrated radiance features and improves weed sensitivity under challenging field variability. The ablation study supports this conclusion: removing the index branch yields a 0.025 mIoU decrease, indicating that index-driven context contributes materially to final predictions. BAWSeg and VISA together offer a reproducible foundation for benchmarking multispectral crop–weed segmentation under realistic spatial and temporal shifts, and for developing models that better generalize across seasons and paddocks. To facilitate follow-up research and practical adoption, we will release the BAWSeg dataset, the VISA implementation, and trained model weights upon publication.
@@
This research received no external funding.
@@
A public preview subset of the underlying multispectral weed-detection dataset is available on IEEE DataPort: Haitian Wang, Xinyu Wang, Muhammad Ibrahim, Dustin Severtson, Ajmal Mian, “Multispectral Remote Sensing for Weed Detection in West Australian Agricultural Lands”, IEEE DataPort, 11 February 2026, doi: https://doi.org/10.21227/f8e1-5934. The full BAWSeg benchmark dataset, together with the VISA implementation source code, trained model weights, and protocol files, will be released upon publication of this article.